OpenStudy (rachie19):

5. A mixture of 2.98 g of H2 and 8.27 g He are placed in a 35.0 L metal container at 410K. a. Calculate the partial pressures of each gas in the mixture. b. Calculate the total pressure in the flask. c. If the flask is heated to 1500K, what is the total pressure inside the flask?

OpenStudy (cuanchi):

a) you need to calculate the number of moles of each sample n=m/MM n=number of moles m=mass MM=molecular mass then with (PV=nRT) calculate the P for each gas P=nRT/V V=35L T=410K R=0.082 atm L/mol K b) total pressure = partial pressure of He + partial pressure of H2 c) P1/T1 = P2/T2 => P2= P1T2/T1

OpenStudy (rachie19):

thank u but can u also help me with some diffusion and effusion problems? @cuanchi

OpenStudy (rachie19):

4. He diffuses 29.45 times faster than an unknown gas (MM He = 4.01 g/mol). Determine the molar mass of the unknown gas.

OpenStudy (rachie19):

6. A compound contains only C, H, and N. It is 58.51 %C, 34.12 %N, and 7.37%H. a. Determine the empirical formula. b. If He effuses 3.2 times as fast as the compound, determine the molecular mass of the compound. c. Using the answer from (b) and (c), determine the molecular formula of the compound.

OpenStudy (cuanchi):

relationship between the ratio of the velocities of the gases and the square root of the ratio of their molar masses